OpenStudy (anonymous):

Find the next three terms of the sequence? -6,-24,-96,-384. A. -768,-1,536,-3,072 B. -768,-3,072, -12,288 C. -1,536,- 6,144,- 24,576 D. - 1,536,- 9,216,- 36,864

OpenStudy (anonymous):

please help ill give medal

OpenStudy (aravindg):

Dont you see any pattern?

OpenStudy (anonymous):

all i knw is that the next number is 768 but i cant find the other @AravindG

OpenStudy (aravindg):

6*4=... 24*4=....

OpenStudy (aravindg):

96*4=....... 384*4=.........

OpenStudy (aravindg):

You got the idea? :)

OpenStudy (anonymous):

yes thank you :)

OpenStudy (anonymous):

its general term is -(6*4n) like putting n=1 then 6 n=2 then 24 n=3 then 96 and so on

OpenStudy (aravindg):

yw :)